SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 8 Apr 2004

Status Report From: NASA HQ
Posted: Thursday, April 8, 2004

image All ISS systems continue to function nominally, except those noted previously or below.

FE Alexander Kaleri conducted Part 2 of the three-part MBI-8 Profilaktika ("countermeasures") fitness test series started yesterday for the fourth time, today using the NS-01 load trainer on the VELO cycle ergometer, supported by tag-up with medical support personnel at TsUP-Moscow. Science Officer Foale was available to assist as necessary. (Last time done: 3/16) [This Russian fitness test consists of four types of exercise, viz., neck tilting (back/forward), simultaneous forearm flexing, trunk extension, and trunk flexes. Each type of exercise consists of a series of 15 motions repeated two times. Load levels are selected by the ground and do not change from test to test. Total duration of the test is 13 min. Gas analysis, subjective evaluation of physical exertion levels, and blood test for lactate and Creatine Kinase levels are also performed as a part of this test, using the TEEM-100M gas analyzer, AccuSport analyzer, and Reflotron-IV blood analyzer.]

With the crew's return to Earth gravity now just three weeks away, Kaleri also had his second orthostatic endurance assessment session with the Russian Chibis suit, conducting the MedOps MO-4 exercise protocol in the below-the-waist reduced-pressure device. CDR Foale assisted his flight engineer as CMO, and the one-hour session was supported by tagup with a ground specialist via VHF. (First time done: 1/13, after 12 weeks in zero-g) [The Chibis ODNT (US: LBNP, lower-body negative pressure) suit provides gravity-simulating stress to the body's cardiovascular/circulatory system for evaluation of Kaleri's orthostatic tolerance (e.g., the Gauer-Henry reflex) after 24 weeks in 0-G.]

CDR/SO Michael Foale began the day with an IMS (inventory management system) and cargo accommodations conference with ground specialists via S-band, addressing stowing issues for the "home stretch" period ahead.

Afterwards, Foale had five hours blocked out on his schedule for reviewing, sorting and consolidating crew-clothing items. [This was part of a recently started IMS activity to search for, audit, consolidate and properly stow equipment to be used by Expedition 9 prior to the arrival of 14P/Progress-249 in May. A crew equipment expert stood by on console to assist Mike with any questions on the clothing dispositions.]

In preparation for the upcoming IFM (in-flight maintenance) to recover the #2 air conditioner (SKV-2) in the Russian segment (RS), which has long been inactive due to a misaligned KVO coolant line (causing an abort of the initial repair attempt on 2/13), the FE used a hacksaw to make cutouts in two SM wall panels (#402, #404) to accommodate the new coolant line and its interference-causing curvature (100mm radius). [This activity was performed today, not yesterday, as erroneously reported on 4/7. The panels are made from 8 mm thick foam material, lined on both sides with a 0.3mm-thick aluminum layer and a noise suppressant mat on one side. To keep resulting FOD (foreign object debris) from floating away, a vacuum cleaner and wet wipes were used, as previously approved by safety specialists, and Mike Foale stood by to assist with FOD avoidance. The treated panels were then reinstalled and photographed for downlink via OCA for ground review. SKV-2 activation is TBD.]

FE Alex Kaleri topped off the water supply of the BIO-5 Rasteniya-2 ("Plants-2") plant growth experiment and completed the regular periodic download of data & imagery collected of the hardware to the Packet Laptop computer for subsequent downlink to the ground via Regul or OCA. [Rasteniya studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.]

Sasha also performed the daily routine technical maintenance of the SM's SOZh environment control & life support system, incl. the toilet equipment (ASU), and prepared the daily IMS (inventory management system) "delta" file for automatic export/import to update the database.

Both crewmembers continued initial preparations for their approaching return to Earth. [Starting yesterday, both have an additional hour per day for coming-home preparations, and their formal workday was accordingly reduced to 5-hrs. 30 min. through 4/20.]

Mike worked out on TVIS treadmill and RED exerciser (Sasha's exercise being part of MBI-8 like yesterday).

At 10:40am EDT, the crew conducted an interactive educational PAO exchange with students at Kirby Middle School in San Antonio, TX, as part of the JSC Middle School Aerospace Scholars (MAS) Program, responding to questions collected from students and uplinked beforehand.

At dinnertime (3:30pm), as every day, the crew supported the Renal Stone prevention experiment by taking the test medication (either potassium citrate or placebo tablets) until the last sample collection phase starting 4/16.

Mike Foale received kudos for his fourth FOOT (Foot/Ground Reaction Forces during Space Flight) exercise ops yesterday, particularly also for the "fantastic" live video. [This concludes his efforts on four excellent sessions during this Increment. The FOOT equipment has been stowed, and the data from the final session, about 8 hours worth, will be downloaded tomorrow.]

The last photo session for BCAT-3 (Binary Colloid Alloy Test-3) is being scheduled for Mike on tomorrow's timeline.

Ground-based troubleshooting is underway on PCU-2 (plasma contactor unit #2) and its leaky latch valve #1 (reported 4/6). [At present, the unit is losing ~1 pound of Xenon gas per week (out of a supply of 68 lbs). Troubleshooting currently consists of repeated attempts at sending Open/Close commands to the valve to seat it properly. Discussions are underway to determine if the recovery can be shortened to avoid loosing additional Xenon. In the worst case, valve #2 would have to be closed, and the PCU would lose redundancy, not availability. [Typical nominal PCU configuration in "Standby" has latch valve #1 open and latch valve #2 closed.]

The onboard laptop "shuffle" continues. Yesterday's attempt by FE Kaleri to restore IBM ThinkPad 760XD laptop "shell" #6060 (with a dimmed display) to take the place of the IP (Internet Protocol) "home phone" machine converted by Mike Foale to a PCS on 4/6, has not met with success, since the 760 "crashes" after ~30 min. of run time, then requires rebooting. [Other steps are being evaluated, and longer-range options are assessed in case of further PCS failure.]

MCC-Houston is preparing for a 4-day test of CMG-3 (control moment gyro #3), designed to evaluate the lubrication situation of its gimbal bearings. The test is scheduled for the period of 4/14-17. [The procedure will involve attitude control handover (ACHO) to the Russian segment (RS) on 4/14, prior to a planned Soyuz thruster test (see below), to allow slewing CMG-3 so as to allow heating of its first gimbal bearing by a near-by electronic box for lubrication evaluation. On 4/16, ACHO to RS will again be required for repositioning the gimbal to heat the other bearing. Thus, during the 4-day period, there will be only two CMGs controlling the station attitude. Since there have been communication problems before between a CMG and the GNC MDM (guidance, navigation & control computer), TDRS accessibility must be maximized during comm coverage.]

Planning is still underway on the ground for further steps at troubleshooting the Lab science window. [Under consideration is another attempt at evacuating the window's inter-pane volume ("Volume D") at a later date, followed by an extended period of pressure checks.]

The Elektron O2 generator continues to operate nominally. Pressure inside the sealed container of the new BZh-7 liquid unit has dropped to 1.08 at, confirming the existence of a tiny leak. When 1.0 at is reached, the unit will again be repressurized with nitrogen (to prevent outside air getting in). [Last time it was repressed: 3/30, which indicates that represses may be required every 10 days. Soyuz 8S will deliver two external micropumps for recovering the operation of a previously failed (and removed) BZh unit. Also, 80 new SFOG (solid-fuel oxygen generator) candles will be manifested on Progress 16P (launch 11/24/04).]

TsUP/Moscow plans for a 7S/Soyuz TMA-3 hot thruster test on 4/14, preparatory to the undocking on 4/29. During the test, the two KDU (integrated propulsion system) manifolds, currently at vacuum, will be independently pressurized with helium gas (He), i.e., first string #2, then string #1. [As reported before, the Soyuz 7S has a very small leak between the He pressurization tanks and the prop tanks of its KDU's manifold #2, which is not perceived as constituting a risk for a nominal descent.]

TsUP also plans to conduct testing of the SM's KURS-P rendezvous radio navigation system on 4/14 in preparation for 8S arrival.

Today's CEO (Crew Earth Observations) targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in "ram"), were Amman, Jordan (ISS had a clear nadir pass over the Jordanian capital city, situated about 30 miles northeast of the Dead Sea), Aral Sea (a series of spring storms continues to sweep across this region of western Asia. However, the crew may have found a good enough break in the weather this pass to document shoreline changes in this rapidly shrinking water body), Sierra Leone Fires (DYNAMIC EVENT: Numerous seasonal fires are indicated by satellite in western Africa, especially Sierra Leone. As ISS approached the African coast from the SW, the crew was to look left of track for oblique views of smoke plumes from these fires), Aegean Sea (as ISS tracked across Greece towards the NE, the crew was to look right of track into the sun glint on the Aegean Sea and try to document internal wave patterns expected to be found there), Baker Island (400mm-lens: Near nadir views in good light and fair weather were expected this pass. Using the long lens to document details of the coral reef structures of this island), Howard Island (400mm-lens: Near nadir views in good light and fair weather were expected this pass. Using the long lens to document details of the coral reef structures of this island), and Lake Winnipeg Ice (DYNAMIC EVENT: As spring slowly comes to the Canadian Prairie Provinces, the crew was to look for the seasonal thawing of the numerous lakes in the region. On this pass trying for indications of the breakup on the ice sheet of the largest ones, Lake Winnipeg and Manitoba, nadir and either side of track).

CEO images can be viewed at the websites.

See also the website "Space Station Challenge" at

U.S. and Russian Segment Status (as of today, 1:14pm EDT).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation TBD). SFOG slot#2 fan suspect (not usable).
  • SM Working Compartment: Pressure (mmHg) -- 742; temperature (deg C) -- 26.2; ppO2 (mmHg) -- 160.4; ppCO2 (mmHg) -- 3.2;
  • SM Transfer Compartment: Pressure (mmHg) -- 761; temperature (deg C) -- 20.4.
  • FGB Cabin: Pressure (mmHg) -- 756; temperature (deg C) -- 23.0;
  • Node: Pressure (mmHg) -- 758.52; temperature (deg C) -- 24.3 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab: Pressure (mmHg) -- 760.37; temperature (deg C) -- 24.9; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Joint Airlock (Equip. Lock): Pressure (mmHg) -- 760.57; temperature (deg C) -- 26.0; shell heater temp (deg C) -- 23.9, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • (n/a = data not available)
  • PMA-1: Shell heater temp (deg C) -- 24.4
  • PMA-2: Shell heater temp (deg C) -- 10.6

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, "sun slicer", i.e., drag reduction-biased by 47 deg).
  • SM batteries: Battery #4 is in "Cycle" mode; all other batteries (7) are in "Partial Charge" mode.
  • FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM); all other batteries (5) are in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3926 kg (8655 lb) as of 3/26/04; [SM(755) + FGB(2512) + Progress M-1(659)]. (Capability: SM -- 860 kg; FGB -- 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- U.S. SIGI-1 (GPS)
  • Angular rate source -- RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 4/14.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings.
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit (as of this morning, 7:03am EDT [= epoch]):

  • Mean altitude -- 364.5 km
  • Apogee -- 371.6 km
  • Perigee -- 357.3 km
  • Period -- 91.83 min.
  • Inclination (to Equator) -- 51.63 deg
  • Eccentricity -- 0.0010584
  • Orbits per 24-hr. day -- 15.68
  • Mean altitude loss last 24 hours -- 150 m
  • Revolutions since FGB/Zarya launch (Nov. '98) -- 30737

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.